How many kilos of tea worth Rs.72 per kg should be mixed with 10kg of tea worth Rs.90 per kg to produce a mixture which will cost Rs.78 per kg
Answer by mananth(16949)

---------------- Price ---------------- quantity
Tea type I 90 ----------------10 pounds
Tea type II 72 ----------------x pounds
Mixture 78 ---------------- 10+x pounds

sum of individual components = quantity in mixture
90*10+72x=78(10 +x)
900+72x =780+78 x
72x -78 x=-120
-6x=-120
x= -120 / -6
x= 20
20 pounds of Tea type II costing Rs. 72 has to be added
m.ananth@hotmail.ca